Hello, I am having trouble computing the amounts for the journal entries that are red. I am using a financial calculator, so I would prefer step by step instructions on how to get the correct answer using a calculator. Thank you!​​​​​​​On January 1, 2020, Sheridan Corporation issued $560,000 of 7% bonds, due in 10 years. The bonds were issued for $601,659, an
Answer #1

Answer -

Journal of Sheridan Corporation

No. Date Account Title and Explanation Debit ($) Credit ($)
(a) Jan. 1, 2020

Cash

Premium on Bonds Payable [Difference]

Bonds Payable

601659

-

-

-

41659

560000

(b) July 1, 2020

Interest Expense [Refer working note - (1)]

Premium on Bonds Payable [Difference]

Cash [($560000 * 7%) * 6/12]

18050

1550

-

-

-

19600

(c) Dec. 31, 2020

Interest Expense [Refer working note - (1)]

Premium on Bonds Payable [Difference]

Interest Payable [($560000 * 7%) * 6/12]

18003

1597

-

-

-

19600

# Working note - (1) - Calculation of Interest Expense -

Date Interest payment ($) Interest Expense ($) Amortization of bond premium ($) Credit balance in bond premium account ($) Credit balance in bonds payable account ($) Book value of the the bonds ($)
(A) (B) (C) = (G) * 3% (D) = (C) - (B) (E) (F) (G) = (F) + (E)
Jan. 1, 2020

19600

[560000 * 3.5%]

- - 41659

560000

[Given in question]

601659

[560000 + 41659]

July 1, 2020

19600

[560000 * 3.5%]

18050

[601659 * 3%]

(1550)

[18050 - 19600]

40109

[41659 - 1550]

560000

[Given in question]

600109

[560000 + 40109]

Dec. 31, 2020

19600

[560000 * 3.5%]

18003​​​​​​​

[600109 * 3%]

(1597)

[18003 - 19600]

38512

[40109 - 1597]

560000

[Given in question]

598512

[560000 + 38512]